OpenStudy (anonymous):

You toss 5 coins. What is the probability that all 5 show heads?

OpenStudy (abb0t):

\((\frac{1}{2})^5\)

OpenStudy (anonymous):



OpenStudy (anonymous):

(1/2)^5=1/32=3.125%

OpenStudy (anonymous):

it works on multiplicative and rule of independent trial of probabilty

OpenStudy (anonymous):

69